Suppose you purchase 516 shares of stock at $29.11 per share with an initial cash investment of $11501. Suppose the call money rate is 3.31 percent and you are charged a 1.5 percent premium over this rate. Calculate your return on investment (in percent) if the share price one year later is $2.50 below the price when you bought the stock. Ignore dividends. Answer to two decimals.

Show steps**
Answer #1

Total purchase = 516*29.11 = 15020.76

Loan = $15020.76- 11501= 3519.76

Interest = 3519.76* (3.31%+1.5%) =169.3

Proceeds from sale = 516* (29.11-2.5) = 13,730.76

Dollar return =13730.76- 11501- 3519.76-169.3 = -1459.30

Return = Dollar return/ Cash outlay

= -1459.30/ 11501

=-12.69%